Workshop ”Analysis of multiple state particle systems”

Date: December 25(Tuesday) - 26(Wednesday), 2001 Place: Keyaki Hall, Chiba Univ. Organizer: Hideki Tanemura (Chiba Univ.)

\underline{\underline{Program}} (Ver. 2001.12.6)

\underline{Dec. 25 (Tue.)}

13:00 - 13:50 Rinaldo Schinazi (Colorado Univ., Univ. of Provence)

On the role of social clusters in the transmission of infectious diseases

14:00 - 14:50 Koji Kuroda (Nihon Univ.)

Application of cluster expansion to a stock price process

15:00 - 15:50 Tomohiro Sasamoto (Tokyo Inst. Tech.)

Particle Condensation for a Two Species Exclusion Process on a Ring

16:00 - 16:50 Yukio Nagahata (Tokyo Inst. Tech.)

Regularity of the diffusion coefficient matrix for the lattice gas with energy in d ≥ 3

\underline{Dec. 26 (Wed.)}

10:00 - 10:50 Rinaldo Schinazi (Colorado Univ., Univ. of Provence） On the spread of drug resistant diseases

11:00 - 11:50 Yasunari Higuchi (Kobe Univ.), Joshin Murai (Okayama Univ.) and J.Wang (Kobe Univ.)

A Dobrushin-Hryniv type CLT for the 2D Widom-Rowlinson model

14:00 - 14:50 Norio Konno (Yokohama National Univ.)

Dualities of stochastic interacting systems

15:00 - 15:50 Tainaka keiichi (Shizuoka Univ.)

Contact process and lattice model for mutualism

16:00 - 16:50 Short comunication

Mail to: tanemura@math.s.chiba-u.ac.jp